OpenStudy (anonymous):

can someone solve for x+y=10 and 10x+15y=115

OpenStudy (johnweldon1993):

x + y = 10 10x + 15y = 115 Lets solve the top equation for 'y' x + y = 10 -x -x ----------- y = -x + 10 Now that we know what y equals...lets plug it into the second equation for every 'y' we see.. 10x + 15y = 115 becomes 10x + 15(-x + 10) = 115 Distribute that 15 into the parenthesis... 10x - 15x + 150 = 115 Combine like terms... -5x + 150 = 115 Subtract 150 from both sides... -5x = -35 Divide both sides by -5 x = 7

OpenStudy (johnweldon1993):

Now that we know x = 7...we can solve for 'y' x + y = 10 right? well if 'x' = 7...y = ? 7 + y = 10 y = 3 So we have x = 7 and y = 3....is that correct...lets check with out 2nd given equation... 10x + 15y = 115 10(7) + 15(3) = 115 70 + 45 = 115 115 = 115 \(\large \color \red {\checkmark} \)